Recent Changes

Friday, February 16

  1. page Trouble Shooting edited ... chapter.tel.stopmoving The Telescope won't move or stopped moving ... the problem. Som…
    ...
    chapter.tel.stopmoving
    The Telescope won't move or stopped moving
    ...
    the problem.
    Sometimes the dome guis get hung up and can cause erratic motion or no motion of the scopes. Check them for current times and continuous updates of numbers. If they are not updating, try to REOPEN them first. If that does not work, close the gui and open a new one. If a new one does not open, the dome server may be dead or Sockman lost track of it. See Dome Server Restart below.
    Azimuth Limit Switches
    ...
    range condition.
    1. On the domegui MANUAL tab, click STOP so pulses won't be sent to the drive by the control software.
    2. Make sure you understand why the limit was hit which may require a trip to the telescope.
    ...
    chapter.tiptilt.clock
    Tiptilt server says the clock isn't running
    ThisFirst check whether the clock itself is running and the other machines receive the clock signal. Look at the clock cards at the back of other computers in the rack. The clock cards have three LEDs, one yellow and two greens. If the computer is receiving the clock signal properly all three LEDs should blink, but at a bug, that I hopedifferent rate. If the LEDs on all the clock cards are solid then reboot the GPS computer. When the GPS computer is down, it is best to fix soon,cycle the power also on the box right above the GPS computer.
    If the clock appears to be working properly on other machines and not on the tiptilt now it is time to reboot tiptilt.
    [There is a bug
    in the
    ...
    the CCD ran. Forran.]
    For
    the time
    ...
    from the lablab. Power OFF the CCD, then reboot the tiptilt machine and go
    ...
    and reboot. Also,When the clock card LEDs in the tiptilt machine indicate proper clock signal, turn the CCD back on and start the tiptilt server.
    Also,
    sometimes Serial
    Sometimes syncing the clock can also cause this problem, but that should be fixed soon. If it does, exit the tiptilt server, log in as root, and reload the tiptilt model using the following commands:
    /sbin/rmmod tiptilt_rt
    (view changes)
    5:00 pm

Monday, January 29

  1. page Classic Guidelines edited ... This is how the power spectrum for good data should look (at 750 Hz sampling, the fringe power…
    ...
    This is how the power spectrum for good data should look (at 750 Hz sampling, the fringe power shows up at 150 Hz):
    {ps_good_2017_12_02_HD26764_005.png}
    ...
    the power spectrum:spectrum (there is a narrow noise peak showing up at the data frequency of 150 Hz, and several other peaks showing up at different frequencies):
    {ps_peaks_2017_12_02_HD26764_004.png}
    (view changes)
    11:50 am
  2. page Classic Guidelines edited ... For stars fainter than K > 5.5 mag, use non-destructive mode if needed. If possible, stay …
    ...
    For stars fainter than K > 5.5 mag, use non-destructive mode if needed.
    If possible, stay at 750 Hz so that the noise peaks are furthest from the fringe window.
    ...
    occurs at a sum of ~ 300
    ...
    the DIFF/SIG viewwindow flatlines at
    You don't have to worry about saturation in destructive mode.
    ...
    determine the best camera speed,
    In H-band you want the sum to be at least 50 counts (H-band has very little background)
    In K-band you want the sum to be at least 50 counts + the background. Close both shutters to estimate the counts in the background.
    In general, use the long scan length. But if the seeing is very good and the fringes aren't bouncing around in the waterfall, you might want to try the short scan.
    Alignment Procedure for Classic
    ...
    before aligning or and before
    ...
    optimize delay BEFOREbefore doing the
    Observing Logs
    Please record the data entries into the electronic observing logs. In the comments section, please make a note of the camera settings and indicate when NIRO alignments are done or if the reference cart position was moved.
    ...
    Other Important Notes
    Keep an eye on the data sequence. Please check to make sure the fringes are in the scan window after the first shutter sequence finishes. This will save time and avoid large gaps in the recording sequence.
    ...
    is ON. If the servo is not on, then the position of the fringes might drift across the window.
    After the fringe data hashave been collected,
    ...
    there are spuriousnoise peaks in
    ...
    data should look:look (at 750 Hz sampling, the fringe power shows up at 150 Hz):
    {ps_good_2017_12_02_HD26764_005.png}
    This is an example of noise spikes from telescope oscillations contaminating the power spectrum:
    (view changes)
    11:44 am
  3. page Classic Guidelines edited ... Keep an eye on the data sequence. Please check to make sure the fringes are in the scan window…
    ...
    Keep an eye on the data sequence. Please check to make sure the fringes are in the scan window after the first shutter sequence finishes. This will save time and avoid large gaps in the recording sequence.
    When clicking SAVE to start recording data, please visually confirm that the SERVO is ON.
    ...
    oscillations? electronic noise?)
    noise?).
    This is how the power spectrum for good data should look:
    {ps_good_2017_12_02_HD26764_005.png}
    This is an example of noise spikes from telescope oscillations contaminating the power spectrum:
    {ps_peaks_2017_12_02_HD26764_004.png}

    (view changes)
    11:34 am
  4. page Classic Guidelines edited ... For stars fainter than K > 5.5 mag, use non-destructive mode if needed. If possible, stay …
    ...
    For stars fainter than K > 5.5 mag, use non-destructive mode if needed.
    If possible, stay at 750 Hz so that the noise peaks are furthest from the fringe window.
    ...
    for saturation whichon the brightest target in the observing sequence. Saturation occurs at
    ...
    per pixel). YouIf the sum of counts is close to 300, then click the "DIFF/SIG" button on the CLASSIC GUI. Saturation occurs when the signal in the DIFF/SIG view flatlines at the end of the scan (maybe we can get a screen shot to show what saturation looks like). If the signal is saturating, then you need to record at a faster (higher) rate.
    You
    don't have
    To determine the camera speed, go to the faintest star in the bracket:
    In H-band you want the sum to be at least 50 counts (H-band has very little background)
    In K-band you want the sum to be at least 50 counts + the background. Close both shutters to estimate the counts in the background.
    In general, use the long scan length. But if the seeing is very good and the fringes aren't bouncing around in the waterfall, you might want to try the short scan.
    Re-alignAlignment Procedure for Classic
    Re-align
    NIRO about
    ...
    aligning or and before moving the reference cart. Both of these can change the calibration.
    When re-aligning, please set a new reference cart position to optimize delay BEFORE doing the NIRO alignment. This is especially important for fast-moving baselines. Ideally, one should wait until the reference cart moves to the correct position before starting the NIRO alignment.
    Observing Logs

    Please record the data entries into the electronic observing logs. In the comments section, please make a note of the camera settings and indicate when NIRO alignments are done or if the reference cart position was moved.
    https://sites.google.com/a/chara-array.org/chara-observing/
    Other Important Notes
    Keep an eye on the data sequence. Please check to make sure the fringes are in the scan window after the first shutter sequence finishes. This will save time and avoid large gaps in the recording sequence.
    When clicking SAVE to start recording data, please visually confirm that the SERVO is ON.
    After the fringe data has been collected, please look at the power spectrum plots to see if there are any problems with noise peaks showing up in the data. If there are spurious peaks in the power spectrum then the cause of these peaks should be investigated (vibrations in the lab? telescope oscillations? electronic noise?)

    (view changes)
    11:31 am

Thursday, January 11

  1. page Operating Procedures edited ... Note: the information contained in these wiki pages needs to be incorporated into the document…
    ...
    Note: the information contained in these wiki pages needs to be incorporated into the documentation in the CVS tree.
    CHARA Array Operating Procedures
    Copyright © 2005-20172005-2018 The CHARA
    ...

    Last updated: 2017-07-212018-01-11
    {Picture_21.png}
    {Picture_22.png}
    (view changes)
    2:35 pm
  2. page Operating Procedures edited ... New S2 Finder Instructions {S2_New_Finder_Camera-r1.docx} Night-time POP Changes Azimuth c…
    ...
    New S2 Finder Instructions {S2_New_Finder_Camera-r1.docx}
    Night-time POP Changes
    Azimuth coudé alignments
    Tip-tilt Splitters Change {Tiptilt Splitters Change.pdf}
    Trouble Shooting
    (view changes)
    2:35 pm
  3. page Azimuth coudé alignments edited There are times when the azimuth coudé alignment needs to be checked and adjusted to establish bet…
    There are times when the azimuth coudé alignment needs to be checked and adjusted to establish better alignment of light to or from a telescope. One symptom of this misalignment is that the tiptilt lock position keeps moving in the ACQ window as the telescope moves with targets across the sky. If you do not have the chance due to time or technical problem to complete the necessary adjustment, please note during observing if a telescope needs the coudé alignment checked, and include this in the technical report.
    To check the coudé alignment, the lab first needs to have the full alignment done for any scope being tested. If you do an alignment during or after observing, make sure the carts are at the back of the rails and that the M10 alignment is good. Also make sure the telescope is being tested on the beam for which it was aligned. When this is done, open the laser and visible shutters, open the M3, M5, and M7 covers, and put in the corner cube for each telescope you will test. Turn on the TV and put it in ACQ mode. Two laser spots should be visible in the ACQ window. Use the Laser Filter to select an appropriate brightness for the spots. Using the Adjust function of the tick marks on the telescope gui, move the marks to the location of the central spot. Use the Zoom function to see the cross hairs over the laser spot you are checking. Set step size to 1 to put the crosshair on the spot as precisely as possible. Write down the position of the crosshairs as displayed in the telescope server.
    Turn on the power to the drives and enable the scopes. Rotate the telescope 180º in increasing azimuth by entering a new SET position (STOW+180º) in the AUTO tab on the Dome gui. Watch the spot position as the scope turns 180º. When it reaches the new position, move the crosshairs to the new position of the laser spot. Note its position in the telescope server as before. Ideally, the spot moves very little or in a circular arc. Sometimes it can move in a noncircular arc, an elliptical arc or even a partial figure 8.
    Calculate the center position of the two points and move the crosshairs to this central position. This will be the new position of the laser spot when the coudé alignment is done. Open an engineering hut gui by typing hutgtk -E S1 & in a terminal window. Substitute each telescope in the example given for the other scopes. In the AOB ENG tab, find the M7_1 and M7_2 entries in the pull down menu next to the [ 100 ] displayed in the white field at the lower left. Select each one and hit the GET button to display the position of the M7 mirror actuators. The values will be centered around 3000000, but can range from 1000000 to 5000000. Write down the value of each actuator as a reference before making any adjustments. If the hut gui cannot connect to the hut server, error messages will be displayed. Try to restart the hut server to reestablish the connection. If that does not work, mention this in a tech report.
    Using the yellow M7 buttons, move the mirror UP, DOWN, LEFT or RIGHT to move the spot to the center of the crosshairs. The default step size is 100000. This may show very small motions and a step size of 200000 can be used if it is moving too slowly. Note that if a large correction is required, for example, the laser spot has moved more than 20-25 pixels between the two positions, a manual adjustment may be needed to center it. The actuators may reach a limit if they are moved too far when trying to make a large correction. Be more cautious if the starting value is already close to one of the limits. The limits are 0 on one end and around 6000000 on the other limit. Once the laser spot has reached the crosshair, write down the new actuator positions for M7_1 and M7_2.
    You may now watch the spot as the telescope is stowed to see if it holds in the new position. Ideally, it will not move, but this will not always be the case. Make sure to close all mirror covers and shutters as needed. Power off each scope when it has stowed. Report your findings and the positions of all spots and actuators to Judit via email when done.
    Version 3, 1-5-18
    Norm

    (view changes)
    2:30 pm

More